# Preconditioners for the Stochastic Training of Implicit Neural Representations

### v1

@@ Abstract
Implicit neural representations have emerged as a powerful technique for encoding complex continuous multidimensional signals as neural networks, enabling a wide range of applications in computer vision, robotics, and geometry. While Adam is commonly used for training due to its stochastic proficiency, it entails lengthy training durations. To address this, we explore alternative optimization techniques for accelerated training without sacrificing accuracy. Traditional second-order optimizers like L-BFGS are suboptimal in stochastic settings, making them unsuitable for large-scale data sets. Instead, we propose stochastic training using curvature-aware diagonal preconditioners, showcasing their effectiveness across various signal modalities such as images, shape reconstruction, and Neural Radiance Fields (NeRF).
footnotetext: * Equal contribution.footnotetext: Correspondence to: Shin-Fang Chng <shinfang.chng@adelaide.edu.au>, Hemanth Saratchandran <hemanth.saratchandran@adelaide.edu.au>
@@ 1 Introduction
Implicit neural representations (INRs) have gained significant traction in the fields of machine learning and computer graphics [57, 60, 12, 28, 10, 56, 35, 55]. Unlike conventional grid-based representations that rely on explicit geometric primitives or parameterized functions [19, 68, 8, 59], INRs take a distinct approach. They define shapes and objects implicitly through trained neural networks. These representations have garnered attention for their ability to efficiently encode complex and highly detailed continuous signals, making them suitable for tasks such as image synthesis [57, 12], shape modeling [35, 56, 11, 38, 1] and robotics [62, 13]. INRs also offer desirable properties, including continuous and differentiable outputs, allowing seamless integration with gradient-based optimizers.
@@
Classically INRs were trained with a ReLU activation, leading to low fidelity reconstructions due to such networks admitting spectral bias [43]. To overcome spectral bias, positional embedding layers (PE) [61] are often added but this increases the parameter count of the network and can lead to noisy gradients. A more recent approach is to use non-traditional activations such as sine [55], Gaussian [44], and wavelets [53]. The use of such activations enable high-frequency encoding without positional embedding layers. The major benefit of these activations over positional embedding layers is their well-behaved gradients [55, 44].
While INRs have opened up new possibilities for realistic and efficient signal representation and manipulation, the training of these architectures still presents a significant research challenge. Curiously, the prevailing approach to training such networks has become almost axiomatic - practitioners typically rely on the Adam optimizer. In the realm of architectures that rely on extensive datasets, the adoption of a stochastic training scheme becomes essential. Surprisingly, the preference for using Adam as the optimizer over a more traditional vanilla optimizer like SGD remains unexplained, despite its widespread usage. In [54], the authors demonstrated remarkable acceleration in training INRs by utilizing second-order optimizers such as L-BFGS [37]. However, an obstacle encountered by [54] was the limited performance of second-order optimizers, such as L-BFGS, in stochastic settings, rendering their methods inefficient for stochastic training of INRs.
In this paper, we investigate the stochastic training of INRs from the perspective of preconditioning. Preconditioners are essential components in optimization and iterative solvers, aiming to enhance convergence and efficiency in numerical methods [50, 37, 5]. They specifically address the condition number, which measures the sensitivity of the solution to small input perturbations. By applying suitable transformations, preconditioners effectively reduce the condition number, facilitating rapid convergence in iterative solvers [16].
We delve into the relationship between the Adam optimizer and preconditioned gradient descent algorithms. Specifically, we demonstrate that Adam can be viewed as employing a diagonal preconditioner based on the Gauss-Newton matrix. This sheds light on why Adam performs remarkably well for INRs compared to vanilla SGD. Furthermore, we propose the use of curvature-aware preconditioners for training INRs, especially for architectures with non-traditional activations such as sine [55], Gaussian [44], or wavelets [53]. We show that employing curvature-aware preconditioners can significantly accelerate training times in stochastic settings.
Our analysis focuses on specific algorithms that leverage these preconditioners, including equilibrated stochastic gradient descent (ESGD) [16], AdaHessian [66], Shampoo [24], and K-FAC [33]. By exploring the potential of these preconditioners, we gain valuable insights into improving the training process for INRs in stochastic settings.
@@
    An analysis of preconditioning for training INRs unveiling the theoretical underpinnings and illuminating the precise scenarios where preconditioning becomes useful.
@@
    An empirical validation of our theory by demonstrating the superiority of curvature-aware preconditioning algorithms, such as ESGD, over Adam across diverse signal modalities, including images, shape reconstruction, and neural radiance fields.
@@ 2 Related Work
Implicit neural representations (INRs) encode signals as weights within a neural network using low-dimensional coordinates, aiming to preserve smoothness in the outputs [68]. The seminal work of Mildenhall et al. [35] demonstrated the power of INRs by showing their remarkable generalization abilities for novel view synthesis. Since this work, there have been several works on INRs finding applications in image view synthesis, shape reconstruction and robotics [14, 17, 20, 36, 39, 40, 41, 45, 51, 55, 15, 63, 67, 52, 69, 9]. Conventionally these works employ a positional embedding layer to their networks, in order to overcome the spectral bias of a ReLU activation [43]. More recently, several new activations have been employed for INRs, such as sine [55], Gaussian [44], wavelets [53]. These new non-traditional activations are showing superior performance on several vision tasks [55, 15, 44, 53, 65, 32].
Preconditioners were first introduced in the context of numerical linear algebra [50, 22], so as to transform ill-conditioned linear systems into better-conditioned forms, enabling faster convergence of iterative methods. Since then they have found their way into machine learning, where they have been widely employed to enhance the efficiency and effectiveness of optimization algorithms. The use of preconditioning techniques in machine learning can be traced back to early works on optimization, such as the conjugate gradient method and its variants [37]. Over time, researchers have explored various preconditioning strategies tailored to the specific characteristics of machine learning problems, including sparse matrices, non-convex objectives, and large-scale datasets [4, 42, 46, 27]. Several well known optimizers such as ADMM [30], mirror descent [2], quasi-Newton methods [7, 21, 37], adaptive optimizers [16, 66, 18, 24, 33] all employ some form of preconditioning.
3 Analysis
3.1 Preconditioners for Stochastic Training
@@
From equation (3) we see that the best preconditioner to use would be the absolute value of the Hessian itself. In the case of a positive definite Hessian, this is precisely what Newton’s method does [37]. The main issue with this preconditioner is that the cost of storing and inverting the preconditioner will be $\mathcal{O}(n^{3})$ for an objective function of $n$ parameters. This will be extremely costly for the training of overparameterized neural networks.
@@
are matrices that represent some form of a diagonal approximation of the Hessian. Examples include the Jacobi preconditioner, which is given by $D^{J}=|diag(H)|$, where ||$\cdot$|| ⋅ | is element-wise absolute value and $H$ is the Hessian. In practical settings, the absolute value is taken so that the preconditioner is positive semidefinite. The equilibration preconditioner is a diagonal matrix which has the 2-norm of each row (or column) on the diagonal: $D^{E}=||H_{i,\cdot}||$. Diagonal preconditioners are often used because of their ability to be stored and inverted in linear time. Diagonal preconditioners are also known as adapative learning rates in the literature.
The algorithm Equilibrated Stochastic Gradient Descent (ESGD) [16] employs $D^{E}$ as a preconditioner to condition stochastic gradient descent. In [16], it is shown that applying ESGD leads to faster training times for autoencoder tasks when compared to vanilla SGD, RMSProp [48] and a preconditioned SGD using $D^{J}$ as preconditioner. The pseudocode for ESGD is given in algorithm 2.
Algorithm 2 ESGD
initial point $x_{0}$, iterations $N$, learning rate $\eta$.
$D\leftarrow 0$
for $t=0,\cdots,N$ do
     $g_{t}\leftarrow\nabla_{x_{t}}f$
     $D_{t}\leftarrow D_{t-1}+D^{E}_{t}$
     $x_{t+1}\leftarrow x_{t}-\eta D_{t-1}^{-\frac{1}{2}}g_{t}$
end for
To efficiently calculate $D^{E}$ during each iteration, it is unnecessary to compute the entire Hessian and then determine the 2-norm of each row, which would consume $\mathcal{O}(n^{2})$ memory for an objective function with $n$ parameters. Alternatively, one leverages the following matrix free estimator result of Bradley and Murray [6]
$||H_{i,\cdot}||=\mathbb{E}[(Hv)^{2}]\text{ for }v\sim\mathcal{N}(0,1).$ (6)
Eqn (6) allows one to compute $D^{J}$ at each iteration with a cost equivalent to one backpropagation.
AdaHessian [66] is a recent algorithm that applies a diagonal preconditioner in the setting of moving averages and thus can be seen as a way of enhancing the Adam optimizer [26]. Given an objective function $f$, AdaHessian, like Adam and RMSProp, computes a first order moment $m_{t}$ and a second order moment $v_{t}$:
$m_{t}$ $=\frac{(1-\beta_{1})\sum_{i=1}^{t}\beta_{1}^{t-i}\textbf{$g_{i}$}}{1-\beta_{1}^{t}}$ (7)
$v_{t}$ $=\bigg{(}\sqrt{\frac{(1-\beta_{2})\sum_{i=1}^{t}D_{i}D_{i}}{1-\beta_{2}^{t}}}\bigg{)}^{k}$ (8)
where $g_{i}$ denotes the gradient iteration $i$, $D_{i}$ denotes the diagonal of the Hessian at iteration $i$ and $0<\beta_{1}$, $\beta_{2}<1$ are hyperparameters that are also used in the Adam optimizer. When $k=1$, equation (8) defines a moving average preconditioner that uses the diagonal of the Hessian, which implies in such cases AdaHessian is using the Jacobi preconditioner $D^{J}$. Algorithm 3 gives the pseudocode for AdaHessian.
Algorithm 3 AdaHessian
initial point $x_{0}$, iterations $N$, learning rate $\eta$, exponential decay rates $\beta_{1}$, $\beta_{2}$, power $k$
$v_{0}\leftarrow 0$
$m_{0}\leftarrow 0$
for $t=1,\cdots,N$ do
     $g_{t}\leftarrow\nabla_{x_{t}}f$
     $D_{t}\leftarrow diag(H_{t})=\text{diagonal of current Hessian}$
     Update $v_{t}$ and $m_{t}$ using eqns (7), (8)
     $x_{t+1}\leftarrow x_{t}-\eta\frac{m_{t}}{v_{t}}$
end for
To efficiently calculate $D^{J}$ during each iteration, one can use the following matrix free estimator result of Bekas et al. [3]
$Diag(H)=\mathbb{E}[v\odot Hv]\text{ for }v\sim\mathcal{R}$ (9)
where $\mathcal{R}$ represents a Rademacher distribution.
@@
Others: There are several second order optimisation algorithms that use preconditioners that do not fall into the above three categories such as, The Gauss-Newton method, The Conjugate Gradient method, Limited memory Broyden-Fisher-Golfarb- Shanno (L-BFGS) [37]. It was shown in [54] that these algorithms do not train well in stochastic settings. Another algorithm of importance for this work is Shampoo, which applies a left and right preconditioning matrix, see [24] for details.
@@ 3.2 Condition Number
In Sec. 3.1, we explained that SGD will move faster through the loss landscape when the curvatures in each direction are similar. Quantitatively the regularity of the loss landscape at a point is measured via the condition number of the Hessian.
Definition 1.
Given a matrix $A$ we define the condition number of A by
$\kappa(A):=\frac{\sigma_{\max}(A)}{\sigma_{\min}(A)}.$ (10)
@@
Using eqn (3) we seek a preconditioner $D$ such that $\kappa((D^{-\frac{1}{2}})^{T}(H(f)(\textbf{x}))(D^{-\frac{1}{2}}))$ is closer to 1 than $\kappa(H)$. This implies that the preconditioner $D$ is regularising the curvatures of the landscape to be equal in each direction.
In [23], Guggenheimer et al. prove the following upper bound on the conditioner number of an $N\times N$ Hessian matrix.
$\kappa(H)<\frac{2}{|det(H)|}\bigg{(}\frac{||H||_{F}}{\sqrt{N}}\bigg{)}.$ (11)
Furthermore, they show that a tight bound is expected when all the singular values, except for the smallest, are roughly equal. In [16], it is shown that by applying an equilibrated preconditioner the above bound can be tightened by a factor of $det(D^{E})\bigg{(}\frac{||H||_{F}}{\sqrt{N}}\bigg{)}^{N}$.
In our empirical validation (based on [16]), we examined a one-dimensional INR scenario. The INR consisted of two hidden layers, each with 16 neurons, aimed at fitting the signal $f(x)=\sin(2\pi x)+\sin(6\pi x)$. During optimization using gradient descent, we employed an equilibrated preconditioner $D^{E}$ and a Jacobi preconditioner $D^{J}$. At each optimization step, we computed both the original Hessian of the gradient descent algorithm and the transformed Hessian using equation (3). Fig. 1 illustrates the experimental results, revealing that the equilibrated descent algorithm exhibits a trajectory with a significantly lower-conditioned Hessian compared to the Jacobi preconditioner and the original Hessian of a gradient descent.
(a) Figure 1: Comparison of condition number (10) before and after applying equilbrated and Jacobi preconditioners on a 1D signal regression. The condition number of the equilbrated preconditioner has significantly reduced (solid blue vs. dotted blue) than Jacobi (solid red vs. dotted red)
(b) Figure 2: Comparison of training convergence (PSNR) between Adam and ESGD on a NeRF task with ReLU-PE. The ReLU-PE INR trained with Adam significantly outperformed the one trained with ESGD, suggesting curvature aware preconditioners offer no benefit for ReLU-PE.
Figure 3: Visualization of the sparsity patterns in the Hessian-vector product matrices for the penultimate layer of ReLU-PE (Left) and Gaussian networks (Right): white pixels represent zero values, and black pixels denotes non-zero values.
3.3 Why INRs train so well with Adam compared to SGD

### v2

# Preconditioners for the Stochastic Training of Neural Fields

@@ Abstract
Neural fields encode continuous multidimensional signals as neural networks, enabling diverse applications in computer vision, robotics, and geometry. While Adam is effective for stochastic optimization, it often requires long training times. To address this, we explore alternative optimization techniques to accelerate training without sacrificing accuracy. Traditional second-order methods like L-BFGS are unsuitable for stochastic settings. We propose a theoretical framework for training neural fields with curvature-aware diagonal preconditioners, demonstrating their effectiveness across tasks such as image reconstruction, shape modeling, and Neural Radiance Fields (NeRF)  [^1] Source code is available at https://github.com/sfchng/preconditioner_neural_fields.git..
Figure 1: We demonstrate that curvature-aware preconditioners can significantly accelerate training convergence for neural fields with suitable activations. Specifically, ESGD (a curvature-aware preconditioned gradient descent algorithm [14]) improves convergence for wavelet, sine, and Gaussian activations, while Adam performs better for ReLU network with positional encoding (ReLU(PE)). In a 2D image reconstruction task (Fig. 1(a)), a wavelet-based neural field trained with ESGD achieves 29.36 PSNR (Fig. 1(e)) at the $10^{\text{th}}$ epoch, outperforming Adam’s 26.61 PSNR (Fig. 1(d)). In a 3D binary occupancy task (Fig. 1(b)), ESGD enables faster convergence for sine activations, as shown by the improved armadillo mesh reconstruction at the $1000^{\text{th}}$ epoch (Fig. 1(g) vs. Fig. 1(f)). In a Neural Radiance Fields (NeRF) task (Fig. 1(c)), a Gaussian-based model trained with ESGD achieves lower reconstruction error (Fig. 1(i), $0.0695$) compared to Adam (Fig. 1(h), $0.0743$), as seen in the error map (darker areas indicate lower error; zoom in 4×4$\times$4 × for clarity).
11footnotetext: Equal contribution. Correspondence to: Shin-Fang Chng $<$shinfangchng@gmail.com$>$, Hemanth Saratchandran $<$hemanth.saratchandran.adelaide.edu.au$>$.
@@ 1 Introduction
In recent years, neural fields, also known as implicit neural representations, have garnered significant attention in machine learning and computer graphics [55, 57, 9, 25, 8, 53, 32, 54]. Unlike traditional grid-based representations [17, 64, 6, 56], neural fields implicitly encode shapes and objects through neural networks. These representations are highly efficient for capturing complex, high-resolution signals, making them well-suited for tasks such as image synthesis [55, 9], shape modeling [32, 53], and robotics [59, 10].
@@
Classically, neural fields trained with ReLU activations suffer from spectral bias and limited fidelity [40]. To address this, positional encoding (PE) layers [58] were introduced but at the cost of increased model complexity and noisy gradients. More recent work has explored non-standard activations, such as sine [54], Gaussian [41] and wavelets [49] which enable high-frequency encoding without positional encodings and produce well-behaved gradients [50].
While neural fields have enabled realistic and efficient representations of continuous signals, progress on minimizing the training time of such representations remains a significant research challenge. Currently, the Adam optimizer is the de facto algorithm for training neural fields, particularly in applications involving datasets where a stochastic training regime is essential. Yet, the preference for Adam over simpler alternatives such as stochastic gradient descent (SGD) has not been rigorously justified for neural fields, despite its widespread adoption. Notably, [50] demonstrated substantial gains in training speed for neural fields using second-order optimizers such as L-BFGS [34]. However, second-order methods are often impractical for stochastic settings [50], limiting their effectiveness for large-scale data, where stochastic training is essential.
In this paper, we approach neural field training from the perspective of preconditioning. Preconditioners improve convergence by reducing the condition number of the optimization problem [46, 34, 4]. We develop a theoretical framework linking Adam’s performance to its implicit use of a diagonal preconditioner derived from the Gauss-Newton matrix and proceed to identify conditions under which curvature-aware diagonal preconditioners, such as those in ESGD (equilibrated stochastic gradient descent) [14], enhance training for neural fields with sine, Gaussian, or wavelet activations. For ReLU activations, we show that such preconditioning provides limited benefit.
Our experiments validate the proposed framework across a range of tasks, including 2D image reconstruction, 3D binary occupancy, and Neural Radiance Fields (NeRF). As demonstrated in Fig. 1, ESGD consistently outperforms Adam for neural fields with wavelet, sine, and Gaussian activations, which, as predicted by our theory, is due to their better-conditioned Hessians. In contrast, Adam performs better for ReLU-PE activations, which are associated with poorly conditioned Hessians. Furthermore, we evaluated other curvature-aware algorithms, including ESGD, AdaHessian [62], Shampoo [21], and K-FAC [30], offering valuable insights into their effectiveness in stochastic training settings.
@@
    We present a theoretical framework for preconditioning in neural field training, uncovering key conditions and highlighting when preconditioning is beneficial.
@@
    We empirically validate our framework, demonstrating the superiority of curvature-aware algorithms like ESGD over Adam across diverse tasks, including image reconstruction, 3D shape modeling, and NeRF.
@@ 2 Related Work
Neural Fields
represent signals as weights within a neural network, using low-dimensional coordinates to maintain smoothness in the output [64]. The influential work by Mildenhall et al. [32] showcased the potential of neural fields for novel view synthesis, demonstrating their impressive generalization capabilities. Since then, neural fields have seen widespread application across image view synthesis, shape reconstruction, and robotics [11, 15, 18, 33, 35, 36, 38, 42, 47, 54, 12, 60, 63, 48, 65, 7]. Traditionally, these approaches integrate a positional embedding layer to counteract the spectral bias introduced by ReLU activations [40]. However, recent work has explored alternative, non-standard activations such as sine [54], Gaussian [41], wavelets [49], and sinc [51]. These activations have shown to outperform conventional approaches in various vision tasks, delivering enhanced performance in capturing high-frequency details and reducing reliance on positional embeddings [54, 12, 41, 49, 61, 29].
Preconditioners
were initially introduced in numerical linear algebra [46, 20] to transform ill-conditioned linear systems into well-conditioned forms, thereby accelerating the convergence of iterative solvers. Over time, preconditioning techniques have become integral to machine learning, where they enhance the efficiency and performance of optimization algorithms. Early applications in machine learning can be traced to foundational optimization methods, such as the conjugate gradient method and its variants [34]. As machine learning problems have grown more complex—often involving sparse matrices, non-convex objectives, and large-scale datasets—researchers have developed specialized preconditioning strategies to meet these unique challenges [3, 39, 43, 24]. Many widely-used optimizers incorporate preconditioning, including methods like ADMM [27], mirror descent [2], quasi-Newton methods [5, 19, 34], and adaptive optimizers [14, 62, 16, 21, 30]. These preconditioned optimizers have become essential for addressing the complex landscapes of modern machine learning models, driving improvements in both convergence speed and solution quality.
3 Notation
The neural fields we consider in this paper will be defined by a depth $L$ neural network with layer widths $\{n_{1},\ldots,n_{L}\}$. We let $X\in\mathbb{R}^{N\times n_{0}}$ denote the training data, with $n_{0}$ being the dimension of the input and $N$ being the number of training samples. The output at layer $k$ will be denoted by $F_{k}$ and is defined by
$F_{k}=\begin{cases}F_{L-1}W_{L}+b_{L},&k=L\\ \phi(F_{k-1}W_{k}+b_{k}),&k\in[L-1]\\ X,&k=0\end{cases}$ (1)
where the weights $W_{k}\in\mathbb{R}^{n_{k-1}\times n_{k}}$ and the biases $b_{k}\in\mathbb{R}^{n_{k}}$ and $\phi$ is an activation applied component wise. The notation $[m]$ is defined by $[m]=\{1,\ldots,m\}$. We will denote the loss function that we train the neural fields with by $\mathcal{L}$. In many experiments in the paper this will be the MSE loss [37]. Our theoretical results will be primarily for the case where $\phi$ is given by one of the activation sine, Gaussian, wavelet or ReLU. For more details on how these activations are used within the context of neural fields we refer the reader to [54, 41, 49].
4 Theoretical Framework
4.1 Overview of Preconditioners
@@
From Eq. 4 we see that the best preconditioner to use would be the absolute value of the Hessian itself. In the case of a positive definite Hessian, this is precisely what Newton’s method does [34]. The main issue with this preconditioner is that the cost of storing and inverting the preconditioner will be $\mathcal{O}(n^{3})$ for an objective function of $n$ parameters [34]. This will be extremely costly for the training of overparameterized neural networks whose Hessians are extremely large.
@@
are matrices that represent some form of a diagonal approximation of the Hessian. Examples include the Jacobi preconditioner, which is given by $D^{J}=|diag(H)|$, where ||$\cdot$|| ⋅ | is element-wise absolute value and $H$ is the Hessian. In practical settings, the absolute value is taken so that the preconditioner is positive semidefinite. The equilibration preconditioner is a diagonal matrix which has the 2-norm of each row (or column) on the diagonal: $D^{E}=||H_{i,\cdot}||$. The preconditioners $D^{J}$ and $D^{E}$ are known as curvature-aware preconditioners as they seek to give a good approximation to the Hessian. Diagonal preconditioners are often used because of their ability to be stored and inverted in linear time. Diagonal preconditioners are also known as adapative learning rates in the literature. The algorithm Equilibrated Stochastic Gradient Descent (ESGD) [14] employs $D^{E}$ as a preconditioner to condition stochastic gradient descent. In [14], it is shown that applying ESGD leads to faster training times for autoencoder tasks when compared to vanilla SGD, RMSProp [45] and a preconditioned SGD using $D^{J}$ as preconditioner. AdaHessian [62] is a recent algorithm that applies a diagonal preconditioner in the setting of moving averages and thus can be seen as a way of enhancing the Adam optimizer [23].
@@
Others: There are several second order optimisation algorithms that use preconditioners that do not fall into the above three categories such as, The Gauss-Newton method, The Conjugate Gradient method, Limited memory Broyden-Fisher-Golfarb- Shanno (L-BFGS) [34]. It was shown in [50] that these algorithms do not train well in stochastic settings.
@@ 4.2 Condition Number
In Sec. 4.1, we explained that SGD will move faster through the loss landscape when the curvatures in each direction are similar. The regularity of the loss landscape at a point is quantified via the condition number of the Hessian.
Definition 4.1.
Given a matrix $A$ we define the condition number $\kappa$ of A by
$\kappa(A):=\frac{\sigma_{\max}(A)}{\sigma_{\min}(A)},$ (7)
@@
where $\sigma$ is the eigenvalue. Using Eq. 4 we observe that if we take a preconditioner $D$ such that $\kappa((D^{-\frac{1}{2}})^{T}(H(f)(\textbf{x}))$ $(D^{-\frac{1}{2}}))$ is closer to 1 than $\kappa(H)$, then this implies that the preconditioner $D$ is regularising the curvatures of the landscape to be equal in each direction, which yields a better training landscape for a gradient descent based algorithm [34].
To empirically illustrate the concept of preconditioning, we examined a one-dimensional curve fitting scenario. Specifically, we considered a feedforward network with two hidden layers, each containing 16 neurons, tasked with fitting the signal $f(x)=\sin(2\pi x)+\sin(6\pi x)$ using a Gaussian activation [41]. During optimization with gradient descent, we applied two types of preconditioners: an equilibrated preconditioner $D^{E}$ and a Jacobi preconditioner $D^{J}$. At each optimization step, we computed both the original Hessian associated with gradient descent and the transformed Hessian using Eq. 4. As shown in Fig. 2, the results demonstrate that applying either the Jacobi or equilibrated preconditioner consistently reduced the condition number of the Hessian compared to the unconditioned case, highlighting the effectiveness of preconditioning in improving optimization.
Figure 2: Fig. 2(a) compares the condition number (7) on a 1D signal regression task before (dotted line) and after (solid line) applying equilbrated $D^{E}$ and Jacobi $D^{J}$ preconditioners. The equilibrated preconditioner significantly reduces the Hessian’s condition number, as shown by the larger gap between the yellow solid and dotted lines compared to the green solid and dotted lines representing the Jacobi preconditioner, thus achieves faster convergence as seen in Fig. 2(b).
4.3 Training with Adam
Figure 3: Comparison of training convergence for neural fields with different activations – ReLU with positional encoding (ReLU (PE)), wavelet, sine and Gaussian – using Adam and SGD optimizers across three tasks: 2D image reconstruction (Fig. 3(a)), 3D binary occupancy reconstruction (Fig. 3(b)), and NeRF (Fig. 3(c)). In all cases, the Adam optimizer outperforms SGD, indicating that leveraging curvature information facilitates faster convergence.
In this section, we aim to explain why training with Adam is generally more efficient than training with SGD for a variety of neural fields. While this observation is well-known within the community, we include it here to provide context and motivation for the theoretical framework developed in Sec. 4.4.